Factory Overhead Rates, Entries, and Account Balance

Sundance Solar Company operates two factories. The company applies factory overhead to jobs on the basis of machine hours in Factory 1 and on the basis of direct labor hours in Factory 2. Estimated factory overhead costs, direct labor hours, and machine hours are as follows:

Factory 1 Factory 2
Estimated factory overhead cost for fiscal
year beginning March 1 $1,018,350 $889,600
Estimated direct labor hours for year 13,900
Estimated machine hours for year 22,630
Actual factory overhead costs for March $81,640 $76,800
Actual direct labor hours for March 1,250
Actual machine hours for March 1,770

a. Determine the factory overhead rate for Factory 1.
$ per machine hour

b. Determine the factory overhead rate for Factory 2.
$ per direct labor hour

Feedback

a-b. Factory 1 overhead is applied based on machine hours. Factory 2 overhead is applied based on direct labor hours.

c. Journalize the entries to apply factory overhead to production in each factory for March.

Factory 1 Work in Process
Factory Overhead
Factory 2 Work in Process
Factory Overhead

Feedback

c. Increase the work in process and reduce the factory overhead.

d. Determine the balances of the factory overhead accounts for each factory as of March 31, and indicate whether the amounts represent overapplied factory overhead or underapplied factory overhead.

Factory 1 $ Debit Underapplied
Factory 2 $ Credit Overapplied
Answer #1

a) Factory overhead rate for Factory 1 = $1,018,350/22630 = $45 per hour

b) Factory overhead rate for Factory 2 = $889,600/13900 = $64 per hour

c) Journal entries for the two factories will be

General Journal Debit Credit
Factory 1 Work in Process 79650
Factory Overhead 79650
Factory 2 Work in Process 80000
Factory Overhead 80000

Factory 1 overhead = $45 * 1770 = $79650

Factory 2 overhead = $64 * 1250 = $80000

d)

For Factory 1:

Applied Overhead = $79650

Actual Overhead = $81640

Under applied overhead = $1990

For Factory 2:

Applied Overhead = $80000

Actual Overhead = $76800

Over applied overhead = $3200

Note - Post any doubts/queries in comments section